how would i write the sum of three consecutive odd numbers is 39 as an equation?
Found 2 solutions by ankor@dixie-net.com, mestrydilip:
Answer by ankor@dixie-net.com(22740)

x + (x+2) + (x+4) = 39
which is
3x + 6 = 39

Answer by mestrydilip(13)

Let us assume the first odd number to be x.
Second number will be x+2.
Third number will be x+4
Sum of the three is given to be 39.
x + (x+2) + (x+4) = 39
x + x+2 + x+4 = 39
3x+6=39